# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v1

@@ 5.2 Taxonomy of Risks
To systematically understand the security landscape in Agentic AI, we categorize risks into four broad classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors.
  • Adversarial Attacks Agents remain vulnerable to prompt injections, gradient-based manipulations, and engineered reasoning traps, risks that are magnified in AMAS due to propagation across agent interactions miao2025autonomous ; liu2025compromising . An illustrative instance is the role-swapping attack observed in ChatDev qian2023chatdev .
  • Data Leakage Persistent memory and extensive inter-agent communication elevate the likelihood of accidental exposure of sensitive information wang2025ip ; mao2025agentsafe . In sensitive domains like financial services and HR, inadequate boundary enforcement and ineffective sanitization amplify these leakage risks tran2025multi ; xu2025forewarned .
  • Agent Collusion and Mode Collapse Coordination mechanisms can inadvertently lead agents to reinforce mutual errors, precipitating groupthink or echo chambers feng2025multi ; cui2025towards . AutoGen experiments illustrate how iterative dialogues among agents can amplify flawed designs, highlighting the risk of emergent misalignment wu2023autogen ; reddy2025enhancing ; cihon2025measuring .
  • Emergent Behavior Complex interactions among agents, memory components, tools, and tasks yield unpredictable behaviors that evade traditional testing and validation methods dawid2025agentic ; la2025large . Agents optimizing for efficiency may unintentionally bypass critical verification steps or suppress contradictory information, scenarios exemplified in blockchain karim2025ai and audio verification contexts li2025audiotrust .
5.3 Case Studies
@@
Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental agentic systems.
@@
Case Study 1: Prompt Leakage in Agentic Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT, where recursive prompt augmentation and insufficient memory controls can lead to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard agentic systems from cascading information leaks agarwal-etal-2024-prompt .
@@
Case Study 2: Collusive Failure in ChatDev. In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of faulty design assumptions. Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations. This scenario underscores the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures qian2024chatdevcommunicativeagentssoftware ; lee2024promptinfectionllmtollmprompt .
@@
Case Study 3: Simulation Attack in Swarm Robotics. In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion. This incident underscores the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences. The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior StrDorFri2024 ; ligot2022using .
@@
Case Study 4: Memory Poisoning in Multi-Agent Chatbots. In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer. This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones. This incident underscores the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities atkins2023memorypoisoning .
@@
Case Study 5: System Prompt Drift in Autonomous Memory Agents In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated. This led to hallucinated goals and emergent behaviors misaligned with initial intentions liu2023lostmiddlelanguagemodels . Points to risks from prompt accumulation and the need for memory version control, audit trails, and reset mechanisms.
These cases illustrate that the introduction of autonomy, memory, and orchestration into LLM-based AI introduces an expanded threat surface that cannot be mitigated with traditional LLM security protocols alone. As agentic systems evolve, new methodologies are needed for rigorous, system-wide threat modeling and runtime assurance that span multiple agents, roles, and memory contexts.
@@ 6 Trust and Explainability in Agentic AI
Agentic AI systems are highly autonomous agents capable of making decisions and taking actions without continuous human oversight. These systems also pose unique challenges and opportunities for human trust. Ensuring that users and stakeholders have confidence in such systems is crucial for their adoption in real-world settings. Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms.
6.1 Building Trust in Autonomous Agents
In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research hughes2025ai . Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable wang2024survey . Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
@@ 6.1 Building Trust in Autonomous Agents
Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent ghafarollahi2024sciagents , MetaGPT ke2025meta Enhances interpretability and debuggability
@@
Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev qian2023chatdev , AutoGen autogen_agents_tutorial Mimics human team transparency; reduces uncertainty
@@
Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow bran2023chemcrow , enterprise dashboards Prevents unsafe execution; reinforces control
@@
Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors cordova2025ai ; strielkowski2025ai Improves predictability; prevents overreach
@@
Social Trust Cues Use polite language, uncertainty expression, and turn-taking GPT-4 safety-tuned profiles kolt2025governing ; kirchhof2025position Enhances user comfort and trust in uncertain scenarios
Transparency and Decision Traceability
@@
Transparency is one of the core enablers of trust. For users to understand and evaluate agentic decisions, the reasoning chains, decision states, and action triggers of agents must be made observable. Several agentic systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent ghafarollahi2024sciagents generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPTke2025meta structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
Status Reporting and Progress Visibility
@@
Clear reporting of intent and intermediate status is also essential. Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions. Tools like AutoGenautogen_agents_tutorial and ChatDevqian2023chatdev have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
Human-in-the-Loop Oversight
@@
Human oversight and intervention mechanisms further reinforce trust. Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority wang2024ali ; kraprayoon2025ai ; sifakis2025reference . Many systems adopt a human-in-the-loop (HITL) paradigm chen2025human , where agents request confirmation before executing high-risk or irreversible actions.
Behavioral Consistency and Bounded Autonomy
@@
Trust requires predictability. Agents should follow defined roles, output in expected formats, and remain within delegated authority. Example In enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow bran2023chemcrow , agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
Social Trust Cues and Language Behaviors
@@
Beyond system-level mechanisms, behavioral consistency and bounded autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols, output style consistency, and language modeling constraints help standardize responses seeber2020collaborating ; acharya2025agentic ; shukla2024ai . Example In adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries cordova2025ai ; strielkowski2025ai . Lastly, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings sandini2024artificial . Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence kolt2025governing ; kirchhof2025position . This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability.
Together, these mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5. As autonomy and complexity increase, combining transparency, oversight, and social alignment will be essential to sustain user confidence.
6.2 Explainability Techniques
Explainability remains a cornerstone in fostering trust, accountability, and reliability in Agentic AI systems, particularly as they operate in high-stakes environments where multi-agent coordination and autonomous decisions directly impact human lives acharya2025agentic . In contrast to conventional AI systems, Agentic AI introduces unique challenges for explainability due to its decentralized architecture, dynamic role assignment, and evolving task decomposition among multiple interacting agents sapkota2025ai .
@@ 6.2 Explainability Techniques
LIME ribeiro2016should Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
SHAP lundberg2017unified Attribution of prediction to input features via Shapley values Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
Decision Provenance Graphs tupayachi2024towards ; bai2024derived Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
Multi-agent SHAP tian2025outlook Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
Symbolic/Rule-Based Agents sapkota2025ai Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
Attention Maps liu2024ava ; hu2024overview Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
Prompt Audit Trails gu2024artificial ; chin2025automating Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
Local Post-hoc Techniques (LIME and SHAP)
Local Interpretable Model-Agnostic Explanations (LIME) ribeiro2016should and SHapley Additive exPlanations (SHAP) lundberg2017unified are widely adopted techniques that offer post-hoc interpretability. LIME approximates a black-box model locally using an interpretable surrogate model, while SHAP attributes predictions to input features via Shapley values. These techniques have been integrated into agentic pipelines, particularly in finance and multi-agent fraud detection systems, where feature-level transparency supports regulatory compliance. However, their direct application in Agentic AI is limited. Each agent may operate with its own objectives, context, and tool access, leading to divergent decision paths that local techniques struggle to reconcile rosenfeld2019explainability ; glanois2024survey .
Explainability in AMAS
Emergent behavior poses another challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system. In platforms such as ChatDev qian2023chatdev or AutoGen wu2023autogen , agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial. To address this, researchers have proposed composite frameworks that combine local explanations with global decision traceability yossef2024explainable ; chander2025toward ; wang2024effects . For example, decision provenance graphs visualize communication flows and interdependencies across agents tupayachi2024towards ; bai2024derived , while causal influence chains track how actions propagate between roles rawal2025causality ; qiu2025reinforcement . Adapted SHAP techniques for multi-agent setups now aim to attribute outcomes to shared memory and agent collaboration tian2025outlook .
Symbolic and Hybrid Architectures
Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures sapkota2025ai . These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding. Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
Lightweight Interpretability: Attention and Prompts
Attention map visualizations have been used to highlight focus areas in multimodal language agents liu2024ava ; ye2024generative ; hu2024overview , offering lightweight but informative views into model behavior. Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction gu2024artificial ; chin2025automating . These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
Ongoing Challenges and Research Directions

### v2

@@ 6.2 Taxonomy of Risks
We propose a taxonomy of risks associated with autonomous multi-agent systems (AMAS), categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 6. These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS. Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
Adversarial Attacks. Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps. These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [120, 121]. For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [17]. Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
Data Leakage. The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information [122, 123]. In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [124, 125]. For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls. Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
Agent Collusion and Mode Collapse. Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [126, 127]. This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies. For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance. Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
Emergent Behaviors. The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [128, 129]. Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability. Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [130], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals [131]. These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
6.3 Real-World Use Cases
@@
Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental Agentic AI systems.
@@
Prompt Leakage in Agentic AI Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [132], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [133]. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks [134].
@@
Collusive Failure in ChatDev. In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of a faulty design [80]. Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations. This scenario highlights the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures [24].
@@
Simulation Attack in Swarm Robotics. In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion [135]. This incident demonstrates the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences. The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior [136].
@@
Memory Poisoning in Multi-Agent Chatbots. In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer. This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones. This incident highlights the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities [137].
@@
System Prompt Drift in Autonomous Memory Agents. In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated. This led to hallucinated goals and emergent behaviors misaligned with initial intentions [138].
@@ 7 Trust and Explainability in Agentic AI
Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent [139], MetaGPT [140] Enhances interpretability and debuggability
@@
Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev [17], AutoGen [108] Mimics human team transparency; reduces uncertainty
@@
Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow [141], enterprise dashboards Prevents unsafe execution; reinforces control
@@
Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors [142, 143] Improves predictability; prevents overreach
@@
Social Trust Cues Use polite language, uncertainty and expressions GPT-4 safety-tuned profiles [144, 145] Enhances user comfort and trust in uncertain scenarios
Agentic AI systems are highly autonomous agents capable of making decisions and taking actions without continuous human oversight. These systems also pose unique challenges and opportunities for human trust. Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms. We show the relationship between trust mechanisms, explainability, and evaluation facets in Figure 7 and discuss next.
7.1 Building Trust in Autonomous Agents
In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment, especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research [146]. Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable [147]. Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
Trust-Enabling Mechanisms Transparency • Status Reporting • HITL • Bounded Autonomy • Social Cues Explainability Techniques LIME • SHAP • Decision Provenance • Attention Maps • Prompt Audits Evaluation Facets Trustworthiness • Explainability • User-Centric • Coordination • Composite
Figure 7: Relationship between trust mechanisms, explainability techniques, and evaluation facets in Agentic AI.
@@ 7.1 Building Trust in Autonomous Agents
Transparency and Decision Traceability. Transparency is one of the core enablers of trust. For users to understand and evaluate agentic decisions, the reasoning chains, decision states, and action triggers of agents must be made observable. Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent [139] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPT[140] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
@@
Status Reporting and Progress Visibility. Clear reporting of intent and intermediate status is also essential. Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions. Tools like AutoGen[108] and ChatDev[17] have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
@@
Human-in-the-Loop Oversight. Human oversight and intervention mechanisms further reinforce trust. Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority [148, 149, 150]. Many systems adopt a human-in-the-loop (HITL) paradigm [151], where agents request confirmation before executing high-risk or irreversible actions.
@@
Behavioral Consistency and Bounded Autonomy. Trust requires predictability. Agents should follow defined roles, output in expected formats, and remain within delegated authority. For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow [141], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
@@
Social Trust Cues and Language Behaviors. Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [152, 16, 153]. For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries [142, 143]. Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [154]. Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [144, 145]. This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability. These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
7.2 Explainability Techniques in AMAS
Emergent behavior of AMAS poses a challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system. In platforms such as ChatDev [17] or AutoGen [18], agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial. While local explainability techniques such as LIME [100] and SHAP [101] remain foundational, their application in Agentic AI is limited due to decentralized agent contexts and shared memory. Researchers have thus proposed composite frameworks that combine local attribution with global decision traceability, such as decision provenance graphs [155], causal influence chains [156], and multi-agent SHAP adaptations [157]. These techniques aim to surface the system-level dynamics that underlie emergent agent behaviors.
Symbolic and Hybrid Architectures. Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures [5]. These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding. Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
Lightweight Interpretability: Attention and Prompts. Attention map visualizations have been used to highlight focus areas in multi-modal language agents [158, 159, 160], offering lightweight but informative views into model behavior. Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction [161, 162]. These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
@@ 7.2 Explainability Techniques in AMAS
LIME [100] Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
SHAP [101] Attribution of prediction to input features via Shapley Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
Decision Provenance Graphs [155, 163] Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
Multi-agent SHAP [157] Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
Symbolic/Rule-Based Agents [5] Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
Attention Maps [158, 160] Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
Prompt Audit Trails [161, 162] Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning